# footnotetext: MSC2020: 53A04, 53A70 (primary), 53A17, 70B15, 37K25, 37K10, 35Q53 (secondary)

### v2

@@ 4.1. A family of open curves with a constant torsion angle
(4.6) $\begin{split}&\kappa_{n}=2\arctan\left(\frac{\widetilde{G}_{n}(0)}{\beta F_{n}(0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv\right)}\right),\end{split}$
where we take the primary value for $\arctan$ in $(-\pi/2,\pi/2)$.
A proof of Theorem 4.1 is given in Appendix B.
4.2. Sufficient conditions for closedness
The constant torsion-angle and unit segment-length curves given in Theorem 4.1 are not necessarily closed. We derive a sufficient condition on the parameters $v,r,y$ for the curve $\gamma$ to be closed to define a Kaleidocycle. We say a curve $\gamma$ is closed with a period $k\in\mathbb{N}$ when $\gamma_{n+k}=\gamma_{n}$ for all $n$. Rewriting this condition using the $\tau$ functions, we obtain
(4.7) $\frac{H_{n+k}}{F_{n+k}}=\frac{H_{n}}{F_{n}},$
It is difficult to give general solutions to (4.7), but we can utilise the pseudo-period of elliptic theta functions (A.12) to obtain particular solutions. For a fixed $k$, choose $m\in\mathbb{Z}\setminus{\{0\}}$ and set $v=my/k$. Then, from (A.12), we see
(4.9) $\Lambda:=\exp\left\{\pi im\left(4r+1\right)\right\}\left(R_{1}R_{3}\right)^{-k}-1=0,$
$\Xi:=\frac{\left(\Delta_{3}+\Delta_{1}\right)k-4\pi im}{\Delta_{3}-\Delta_{1}}=0.$
Note that since
(4.10) $\frac{\partial}{\partial r}\log(\Lambda+1)=4\pi im-k\left(\Delta_{3}+\Delta_{1}\right)=-\left(\Delta_{3}-\Delta_{1}\right)\Xi$
holds, (4.9) can also be written as
(4.11) $\Lambda=0,\quad\frac{\partial}{\partial r}\Lambda=0.$
From the above discussion we obtain the following theorem.
@@ 4.2. Sufficient conditions for closedness
For a fixed $k$, let $m\in\mathbb{Z}\setminus{\{0\}}$, $y>0$, $r\in\mathbb{R}\setminus{(1/2\mathbb{Z})}$ and $v\in\mathbb{R}\setminus{(y\mathbb{Z})}$ satisfy $kv=my$ and (4.9). Then, the curve defined by Theorem 4.1 is closed and forms a $k$-Kaleidocycle with a torsion angle given by (4.5).
@@
The closure condition (4.7) holds for any $n\in\mathbb{Z}$ when the parameters $v,r,y$ satisfy $kv=my$ and (4.9). Thus, the conditions $kv=my$ and (4.9) give sufficient conditions for the curve given by Theorem 4.1 to be a closed curve. ∎
Hereafter, we assume that the parameters $v,r,y$ are taken to satisfy the sufficient conditions for closedness.
Example 4.3.
The classical 6-Kaleidocycle (Figure 1 Left) with $\lambda=\pi/2$ (e.g., see [8]) is obtained by $(v,r,y,m)\approx(0.640,0.250,1.279,3)$.
@@
By using (4.9) and (A.12), we see that $\widetilde{B}_{0}=(-1)^{m}\widetilde{B}_{k}$. From this, we also see that $m$ and $Lk$ have the same parity. Furthermore, under the transformation $\widetilde{B}_{n}\rightarrow(-1)^{n+1}\widetilde{B}_{n}$, $\widetilde{N}_{n}\rightarrow(-1)^{n+1}\widetilde{N}_{n}$, the signed curvature angle becomes $(-1)^{n+1}\kappa_{n}$ and the torsion angle becomes $\pi+\lambda$. This indicates that this transformation corresponds to the one described in Proposition 5.16.
5. Construction of the motion of a Kaleidocycle
By introducing the deformation parameter to the curve constructed in §4.1, we construct an explicit formula for the motion of a Kaleidocycle in Theorem 5.1. We then derive the evolution equations for the vertex positions of the curve and the (potential of the) curvature. Connections to the previously-known isoperimetric deformation of a curve by the semi-discrete mKdV equation [16] and by the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation [19] are discussed and summarised in Table 2.
5.1. Explicit formula of the discrete curve with deformation parameter
We introduce the deformation parameter so that only the signed curvature angle changes with a constant torsion angle and while satisfying (3.1)-(3.6). Additionally, the real parameters $C$ and $\Gamma$ are introduced to control the global motion of the curve. The following theorem describes a Kaleidocycle together with its motion explicitly in terms of elliptic theta functions.
Theorem 5.1.
Denote $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$. Consider the following functions parameterised by $t\in\mathbb{R}$.
(5.1) $F_{n}(t,z)=\alpha_{2}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
$f_{n}(t,z)=\vartheta_{3}\left(-\frac{1}{2}iv+r\right)R_{3}^{n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz-\frac{\Gamma}{2}it\right)\vartheta_{2}\left(\mu_{n}+r+it\right),$
$g_{n}(t,z)=\vartheta_{1}\left(\frac{1}{2}iv+r\right)R_{1}^{-n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz+\frac{\Gamma}{2}it\right)\vartheta_{4}\left(\mu_{n}-r+it\right),$
$G_{n}(t,z)=\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}R_{1}^{-n}R_{3}^{n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
$H_{n}(t,z)=\frac{u_{2}}{\alpha_{2}s_{2}}R_{1}^{-n}R_{3}^{-n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz+\Gamma it\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it\right),$
$\widetilde{G}_{n}(t,z)=-i\alpha_{2}\frac{\vartheta_{1}(iv)}{\vartheta_{3}(0)}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
where $C,\Gamma\in\mathbb{R}$ are constants. Then, they satisfy (3.1)-(3.6) to form $\tau$ functions. For each $t$, the corresponding discrete curve $\gamma$ (3.10) has a unit segment length $\left|\gamma_{n+1}-\gamma_{n}\right|=1$. The torsion angle is given by
(5.2) $\lambda_{n}=\frac{1}{i}\log\left(\frac{G_{n}^{*}(t,0)G_{n-1}(t,0)}{\widetilde{G}_{n}(t,0)\widetilde{G}_{n-1}(t,0)}\right)=\frac{1}{i}\log\left(R_{1}R_{3}^{-1}\right),$
(5.3) $\begin{split}&\kappa_{n}\left(t\right)=2\arctan\left(\frac{\widetilde{G}_{n}(t,0)}{\beta F_{n}(t,0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv+it\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv+it\right)}\right),\end{split}$
where we take the primary value for $\arctan$ in $(-\pi/2,\pi/2)$. If $r,v,y$ are chosen as in Theorem 4.2, $\gamma$ is closed for every $t$.
A proof of Theorem 5.1 is given in Appendix B.
Proposition 5.2.
@@ 5.1. Explicit formula of the discrete curve with deformation parameter
Proof.
For $A\in\mathbb{C}$ and $B:\mathbb{C}\to\mathbb{C}$, we have
(5.4) $\begin{split}&\left(\begin{array}[]{ccc}\cos\Gamma t&-\sin\Gamma t\\ \sin\Gamma t&\cos\Gamma t\\ \end{array}\right)\left(\begin{array}[]{ccc}A+A^{*}\\ \frac{1}{i}\left(A-A^{*}\right)\end{array}\right)\\ &=\frac{1}{2}\left(\begin{array}[]{ccc}e^{i\Gamma t}+e^{-i\Gamma t}&-\frac{1}{i}\left(e^{i\Gamma t}-e^{-i\Gamma t}\right)\\ \frac{1}{i}\left(e^{i\Gamma t}-e^{-i\Gamma t}\right)&e^{i\Gamma t}+e^{-i\Gamma t}\\ \end{array}\right)\left(\begin{array}[]{ccc}A+A^{*}\\ \frac{1}{i}\left(A-A^{*}\right)\end{array}\right)\\ &=\left(\begin{array}[]{ccc}e^{i\Gamma t}A+e^{-i\Gamma t}A^{*}\\ \frac{1}{i}\left(e^{i\Gamma t}A-e^{-i\Gamma t}A^{*}\right)\end{array}\right)\end{split}$
and
(5.5) $\begin{split}2\frac{\partial\log\left(e^{\frac{C}{2}tz}B\right)}{\partial z}=Ct+2\frac{\partial\log B}{\partial z}.\end{split}$
The assertion follows from the definition (3.10). ∎
@@
Table 1 summarises the relationship between various variables. The first column lists the parameters that determine the curve $\gamma$ through the variables listed in the first row. A circle in the table indicates that the variable shown in the first row is dependent on the parameter shown in the first column.
Table 1. Dependency of variables on the parameters $v,r,y,k,m,t,C,\Gamma$.
$\gamma$ $\kappa$ $\lambda$ $\Lambda$ $\Xi$ $u_{2}$ $\beta$ $s_{2}$ $R_{j}$ $\Delta_{j}$ $\alpha_{2}$ $\beta_{1,2}$ $\epsilon_{1,2}$ $\omega$ $\delta_{1,2,3}$
$v$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$r$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$y$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$k$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$m$circ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$ $\circ$
$t$circ $\circ$
$\Gamma$circ $\circ$
$C$circ
(3.10) (5.3) (5.2) (4.9) (4.9) (4.2) (4.2) (4.2) (4.2) (4.2) (4.2) (5.6) (5.6) (5.6) (5.6)
The constructed motion has a periodic curvature. The next proposition follows from (A.12).
Proposition 5.3.
We have $\kappa_{n}\left(t+y\right)=-\kappa_{n}\left(t\right)$. In other words, $2y$ is a period of the motion.
Remark 5.4.
There are certain relations between the curves constructed in Theorem 5.1 with different sets of parameters. The transformation of parameters $(v,m)\rightarrow(-v,-m)$ preserves the closure condition. We see that $\kappa_{n}\left(t;-v\right)=-\kappa_{n}\left(-t;v\right),\lambda(-v,r)=-\lambda(v,r)$. In what follows, We always choose $(v,m)$ or $(-v,-m)$ so that $\lambda\geq 0$.

### v3

@@ 4.1. A family of curves with a constant torsion angle
(4.13) $\begin{split}\kappa_{n}&=2\arctan\left(\frac{\widetilde{G}_{n}(0)}{\beta_{1}F_{n}(0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv\right)}\right).\end{split}$
This theorem follows as a corollary of Theorem 5.1 and will be proven in §5.1.
Remark 4.3.
Note that $2\tan\frac{\kappa_{n}}{2}$ is often considered as the discritised curvature (see, for example, [14]), which is expressed using Jacobi $\mathrm{cn}$-function as follows:
[…]
where $\tilde{K}=\pi\vartheta_{3}\left(0;\frac{i}{y}\right)^{2}/2$ is the complete elliptic integral of the first kind and $\tilde{\xi}=\vartheta_{2}\left(0;\frac{i}{y}\right)^{2}/\vartheta_{3}\left(0;\frac{i}{y}\right)^{2}$ is the elliptic modulus.
5. Deformation of a discrete curve with a constant torsion angle
By introducing the deformation parameter $t$ to the curve constructed in §4.1, we construct an explicit formula for the motion of a Kaleidocycle in Theorem $5.1$). We then derive the evolution equations for the vertex positions of the curve and the (potential of the) curvature (§5.2, §5.3, §5.4). Connections to the previously-known isoperimetric deformation of a curve by the semi-discrete mKdV equation [18] and by the semi-discrete potential mKdV equation and the semi-discrete sine-Gordon equation [21] are discussed. This is also summarised in Table 1 in §6.
5.1. Explicit formula of the curve with deformation parameter
We introduce the deformation parameter $t$ so that only the signed curvature angles change while the torsion angle is kept constant. Additionally, the real parameters $C$ and $\Gamma$ are introduced to control the global motion of the curve. The following theorem describes a Kaleidocycle together with its motion explicitly in terms of elliptic theta functions.
@@ 5.1. Explicit formula of the curve with deformation parameter
Denote $\mu_{n}=ivn+\frac{z}{\Delta_{3}-\Delta_{1}}$. Consider the following functions parameterised by $t\in\mathbb{R}$:
(5.1) $F_{n}(t,z)=\alpha_{1}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
$f_{n}(t,z)=\alpha_{2}R_{3}^{n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz-\frac{\Gamma}{2}it\right)\vartheta_{2}\left(\mu_{n}+r+it\right),$
$g_{n}(t,z)=\alpha_{3}R_{1}^{-n}\exp\left(\frac{\left(n+\frac{1}{2}\right)\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz+\frac{\Gamma}{2}it\right)\vartheta_{4}\left(\mu_{n}-r+it\right),$
$G_{n}(t,z)=\alpha_{4}R_{1}^{-n}R_{3}^{n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
$H_{n}(t,z)=\alpha_{5}R_{1}^{-n}R_{3}^{-n}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz+\Gamma it\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it\right),$
$\widetilde{G}_{n}(t,z)=-i\alpha_{4}\exp\left(\frac{n\Delta_{3}}{\Delta_{3}-\Delta_{1}}z+\frac{C}{2}tz\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv+it\right),$
where $C,\Gamma\in\mathbb{R}$ are constants. Then, they satisfy (3.1)-(3.6) with $\beta_{1}$ given by (4.11) and form $\tau$ functions. For each $t$, the corresponding discrete curve $\gamma$ (3.10) has a unit segment length $\left|\gamma_{n+1}-\gamma_{n}\right|=1$. The torsion angle is given by
(5.2) $\begin{split}\lambda_{n}&=\arccos\left(\frac{G_{n}^{*}(t,0)G_{n-1}(t,0)+G_{n}(t,0)G_{n-1}^{*}(t,0)}{2\widetilde{G}_{n}(t,0)\widetilde{G}_{n-1}(t,0)}\right)\\ &=\arccos\left(\frac{R_{1}R_{3}^{-1}+R_{1}^{-1}R_{3}}{2}\right).\end{split}$
(5.3) $\begin{split}\kappa_{n}\left(t\right)&=\kappa_{n}\left(t;v,y\right)=2\arctan\left(\frac{\widetilde{G}_{n}(t,0)}{\beta_{1}F_{n}(t,0)}\right)\\ &=2\arctan\left(-i\frac{\vartheta_{1}(iv)\vartheta_{4}\left(ivn-\frac{1}{2}iv+it\right)}{\vartheta_{3}(iv)\vartheta_{2}\left(ivn-\frac{1}{2}iv+it\right)}\right).\end{split}$
@@
Putting
[…]
in (A.2) yields (3.1). Putting
[…]
in (A.2) yields (3.2). Putting
[…]
in (A.3) yields (3.3). Putting
[…]
in (A.4) yields
(5.24) $\begin{split}&\frac{\vartheta_{2}\left(0\right)\vartheta_{4}\left(0\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\vartheta_{4}\left(\mu_{n+1}-\frac{1}{2}iv-2r+it\right)\\ &=\frac{\vartheta_{2}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(-\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(\mu_{n}-r+it\right)\vartheta_{4}\left(\mu_{n}-r+it\right)\\ &-\vartheta_{3}\left(\mu_{n}-r+it\right)\vartheta_{1}\left(\mu_{n}-r+it\right).\end{split}$
Also, putting
[…]
in (A.4) yields
(5.30) $\begin{split}&\frac{\vartheta_{2}\left(0\right)\vartheta_{4}\left(0\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it\right)\\ &=\frac{\vartheta_{2}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}\vartheta_{2}\left(\mu_{n}-r+it\right)\vartheta_{4}\left(\mu_{n}-r+it\right)\\ &-\vartheta_{3}\left(\mu_{n}-r+it\right)\vartheta_{1}\left(\mu_{n}-r+it\right).\end{split}$
Subtracting (5.30) from (5.24) and substituting
(5.31) $\begin{split}&\vartheta_{3}\left(2r\right)\vartheta_{4}\left(0\right)\vartheta_{2}\left(0\right)\vartheta_{1}\left(iv\right)\\ &=\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(-\frac{1}{2}iv+r\right)\\ &-\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{2}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\frac{1}{2}iv+r\right),\end{split}$
in the equation, we obtain
(5.32) $\begin{split}&\frac{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{3}(2r)\vartheta_{1}(iv)}\\ &\times\left\{\frac{\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)}{\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(-\frac{1}{2}iv+r\right)}\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\vartheta_{4}\left(\mu_{n+1}-\frac{1}{2}iv-2r+it\right)\right.\\ &\left.-\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\vartheta_{4}\left(\mu_{n}-\frac{1}{2}iv-2r+it\right)\right\}\\ &=\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(\mu_{n}-r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\mu_{n}-r+it\right),\end{split}$
from which we see (3.4) holds. Putting
[…]
in (A.2) yields
(5.38) $\begin{split}&\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it+h\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it-h\right)\vartheta_{3}(2r)\vartheta_{3}(0)\\ &=\vartheta_{2}\left(\mu_{n}+r+it\right)\vartheta_{2}\left(\mu_{n}-r+it\right)\vartheta_{3}\left(\frac{1}{2}iv+r+h\right)\vartheta_{3}\left(\frac{1}{2}iv-r+h\right)\\ &-\vartheta_{4}\left(\mu_{n}+r+it\right)\vartheta_{4}\left(\mu_{n}-r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r+h\right)\vartheta_{1}\left(\frac{1}{2}iv-r+h\right).\end{split}$
Differentiating (5.38) with $h$ and then putting $h=0$, we obtain
(5.39) $\begin{split}&\vartheta_{3}(2r)\vartheta_{3}(0)\left(\Delta_{3}-\Delta_{1}\right)D_{z}\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\cdot\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\\ &=-\Delta_{3}\vartheta_{3}\left(-\frac{1}{2}iv+r\right)\vartheta_{3}\left(\frac{1}{2}iv+r\right)\vartheta_{2}\left(\mu_{n}+r+it\right)\vartheta_{2}\left(\mu_{n}-r+it\right)\\ &-\Delta_{1}\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\mu_{n}+r+it\right)\vartheta_{4}\left(\mu_{n}-r+it\right).\end{split}$
Using the equation (5.38) with $h=0$ and substituting (5.39), we obtain
(5.40) $\begin{split}&\vartheta_{3}(2r)\vartheta_{3}(0)\left\{\left(\Delta_{3}-\Delta_{1}\right)D_{z}\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\cdot\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\right.\\ &\left.+\Delta_{3}\vartheta_{2}\left(\mu_{n+1}-\frac{1}{2}iv+it\right)\vartheta_{2}\left(\mu_{n}-\frac{1}{2}iv+it\right)\right\}\\ &=\left(\Delta_{3}-\Delta_{1}\right)\vartheta_{1}\left(-\frac{1}{2}iv+r\right)\vartheta_{4}\left(\mu_{n}+r+it\right)\vartheta_{1}\left(\frac{1}{2}iv+r\right)\vartheta_{4}\left(\mu_{n}-r+it\right).\end{split}$
From this we see (3.5) holds. By the definition of $G$ and $\widetilde{G}$, (3.6) follows. Theorem 4.2 follows from Theorem 5.1 by setting $t=0$. ∎